A pelican is sitting on its perch in the habit as ocean waves come splashing in. The crests is the waves drench the pelican as it collided with its perch. The drenching occurs every 1.40 seconds. The crest are positioned 9.8 meters apart. What is the period of the wave? What is the frequency of the wave? Determine the speed of the Wave?
1
Expert's answer
2020-04-27T10:14:56-0400

"T=1.40\ s"

"f=T^{-1}=1.4^{-1}=0.714\ Hz"

"v=\frac{\lambda}{T}=\frac{9.8}{1.4}=7\frac{m}{s}"